OpenStudy (anonymous):

-2x+15y+z=44 4x+3y+3z=18 -3x+6y-z=18

OpenStudy (anonymous):

What do you have to the problem?

OpenStudy (cadteach):

For ease of communicating, lets label your equations A, B, and C, in the order you wrote them. So if I say A, I'm talking about -2x+15y+z =44, OK? We will solve for x,y,and z, by using elimination, 2 at a time. If I add equations A and C you will see that the +z in A and the -z in C are additive inverses, so they add to zero. SO , then adding A and C looks like -2x+15y+z =44 + -3x +6y - z=18 --------------------------- -5x + 21y = 62 Next, I will multiply 3 time A, so -6x + 45y +3z = 132 (from 44*3) Then, subtracting 3*A - B gives -3x + 39 y = 114 (from 132 -18) Now, perform elimination on -5x +21y = 62 and -3x +39y = 114 (we'll use 3 times the first, and -5 times the second) . Does this help at all?